# A Personalised 3D+t Mesh Generative Model for Unveiling Normal Heart Dynamics

### v1

@@ Abstract
Understanding the structure and motion of the heart is crucial for diagnosing and managing cardiovascular diseases, the leading cause of global death. There is wide variation in cardiac shape and motion patterns, that are influenced by demographic, anthropometric and disease factors. Unravelling the normal patterns of shape and motion, as well as understanding how each individual deviates from the norm, would facilitate accurate diagnosis and personalised treatment strategies. To this end, we developed a novel conditional generative model, MeshHeart, to learn the distribution of cardiac shape and motion patterns. MeshHeart is capable of generating 3D+t cardiac mesh sequences, taking into account clinical factors such as age, sex, weight and height. To model the high-dimensional and complex spatio-temporal mesh data, MeshHeart employs a geometric encoder to represent cardiac meshes in a latent space, followed by a temporal Transformer to model the motion dynamics of latent representations. Based on MeshHeart, we investigate the latent space of 3D+t cardiac mesh sequences and propose a novel distance metric termed latent delta, which quantifies the deviation of a real heart from its personalised normative pattern in the latent space. In experiments using a large dataset of 38,309 subjects, MeshHeart demonstrates a high performance in cardiac mesh sequence reconstruction and generation. Features defined in the latent space are highly discriminative for cardiac disease classification, whereas the latent delta exhibits strong correlation with clinical phenotypes in phenome-wide association studies. The codes and models of this study will be released to benefit further research on digital heart modelling.
@@ Introduction
The heart is one of the most important and vital organs within the human body [1]. It is composed of four morphologically distinct chambers that function in a coordinated manner. The shape of the heart is governed by genetic [2] and environmental factors, as well as a remodelling process observed in response to myocardial infarction, pressure overload, and cardiac diseases [2, 3]. The motion of the heart follows a periodic non-linear pattern modulated by the underlying molecular, electrophysiological, and biophysical processes [4]. The unveiling of complex patterns of cardiac shape and motion will provide important information to assess the status and function of the heart for both clinical diagnosis and cardiovascular research [5, 6, 7, 8].
@@
The current state-of-the-art for assessing cardiac shape and function is to perform analyses of cardiac images, e.g., cardiac magnetic resonance (MR) images, and extract imaging-derived phenotypes (IDPs) of cardiac chambers [bai2020population,Dewey2020]. Most imaging phenotypes, such as chamber volumes or ejection fractions, provide a global and simplistic measure of the complex 3D-temporal (3D+t) geometry of cardiac chambers [9]. These measures may not fully capture the dynamics and variations of heart function across individuals. Recognising these limitations underscores the importance of establishing a more precise model of cardiac status, defining what a normal heart looks like and moves like. Nevertheless, it is a non-trivial task to describe the normative pattern of the 3D shape or even 3D-t motion of heart, due to the complexity in representing high-dimensional spatio-temporal data.
@@
Recently, machine learning techniques have received increasing attention for cardiac shape and motion analysis [4, 10, 11]. Most existing research focuses on developing discriminative machine learning models, that is, training a model to perform classification tasks between different shapes or motion patterns [12, 13, 4, 6]. However, discriminative models offer only classification results and do not explicitly explain what the normative pattern of cardiac shape or motion looks like [14]. In contrast, generative machine learning models provide an alternative route. Generative models are capable of describing distributions of high-dimensional data, such as images [15, 16, 17], geometric shapes [18, 19, 20], or molecules [21, 22], which allow the representation of normative data patterns in the latent space of the model. Although generative models have been explored for cardiac shape modelling [23, 24, 25, 26, 27] or for cardiac imaging data augmentation [28], their application to personalised normative modelling of the heart has been less explored.
@@
Figure 1: An overview of the MeshHeart model. a, Model architecture: MeshHeart encodes a sequence of cardiac meshes using a mesh encoder $M_{enc}$ and encodes clinical factors using a conditional encoder $C_{enc}$. The encoder outputs across the time frames and along with distribution tokens $\mu_{token}$ and $\Sigma_{token}$, are processed by a temporal Transformer encoder $T_{enc}$. A Transformer decoder $T_{dec}$ and a mesh decoder $M_{dec}$ generate a 3D cardiac mesh sequence based on clinical conditions. b, Given a set of clinical factors, an example of the generated mesh sequence across time frames. c, Conceptual framework: MeshHeart constructs a normative cardiac mesh sequence using personal information including age, sex, weight, and height. A real heart can be compared to its personalised norm by the latent vector. The latent delta $\Delta z$ is the computed distance between the latent vector of a synthetic normal heart (blue dot) and that of the real heart (red dot).
@@ Results
Utilising the latent representations learnt by MeshHeart, we assessed the ability of the model to generate new synthetic cardiac mesh sequences that mimic real heart dynamics. To evaluate the fidelity and diversity of the generation, we calculated the similarity between the distributions of real meshes and generated synthetic meshes. For each real heart in the test set ($n=4,000$), we applied MeshHeart to generate synthetic mesh sequences using the same clinical factors (age, sex, weight, and height) as the individual as the model input. During the generation stage, we chose 20 random samples from the Gaussian distribution of the latent space and generated the corresponding mesh sequences. For both real and synthetic meshes, clinically relevant metrics for cardiac structure and function were derived, including left ventricular end-diastolic volume (LVEDV), LV end-systolic volume (LVESV), LV ejection fraction (LVEF), LV myocardial mass (LVM), right ventricular end-diastolic volume (RVEDV), RV end-systolic volume (RVESV) and RV ejection fraction (RVEF). For each metric $m$, its probability distributions against age $P(m|c=age)$ and against sex $P(m|c=gender)$, were calculated. The similarity between real and synthetic probability distributions was quantified using the Kullback-Leibler (KL) divergence [32] and the Wasserstein distance (WD) [33], with a lower value denoting a higher similarity, i.e. better generation performance. KL divergence is a metric from information theory that evaluates the dissimilarity between two probability mass functions. Similarly, WD measures the dissimilarity between two probability distributions. MeshHeart’s ability to replicate real data distributions is quantitatively demonstrated in Figure 3(a). MeshHeart achieves lower KL and WD scores compared to existing methods, as shown by radar plots with the smallest area, suggesting that the synthetic data generated by the proposed model closely aligns with the real data distribution for clinically relevant metrics. Supplementary Tables S3 and S4 report the detailed KL divergence and WD scores for different methods.
@@
We also examined the latent representation learnt by MeshHeart using t-SNE (t-distributed Stochastic Neighbour Embedding) visualisation [34] as illustrated in Supplementary Figure S1. The t-SNE plot projects the 64-dimensional latent representation of a mesh, extracted from the last hidden layer of the Transformer Encoder $T_{enc}$, onto a two-dimensional space, with each point denoting a mesh. It shows 10 sample sequences. For each sample, the latent representations of the meshes across time frames form a circular pattern that resembles the rhythmic beating of the heart [35].
Figure 4: Comparison of disease classification performance, in terms of the AUC (area under the curve) scores, when different feature sets are used. The three feature sets being compared include "Phenotypes $+$ Confounders (Age, Sex)", "Latent vector $+$ Confounders", and "Phenotypes $+$ Latent vector $+$ Confounders", with each feature set represented by a unique color in the plot. The three subplots show the performance of three different classifiers, AdaBoost, LDA, and SVM. The x-axis denotes the disease type and the y-axis denotes the AUC score.
@@
After demonstrating the generative capability of MeshHeart, we explore its potential for clinical applications, in particular using its latent space which provides a low-dimensional representation of cardiac shape and motion. The latent feature analyses were conducted on 17,309 subjects. More than half (58.5%) had a reported diagnosis of at least one disease. We employ the latent vector $z$ of each mesh sequence, a 64-dimensional vector, as the feature for correlation analysis and for cardiac disease classification. Figure 5(a) shows that the latent vector exhibits strong correlations with conventional imaging phenotypes, such as LVM, LVEDV, and RVEDV etc. Figure 4 and Supplementary Table S6 compare the classification performance of six cardiac diseases when using different feature sets. The three evaluated feature sets include "Phenotypes + Confounders (Age, Sex)", "Latent Vector + Confounders", and "Phenotypes + Latent Vector + Confounders". The classification performance is evaluated using the AUC (area under the curve) scores for three different classifiers: AdaBoost, linear discriminant analysis (LDA), and support vector machine (SVM). The six cardiovascular diseases include myocardial infarction (ICD-10 code I21), ischemic heart diseases (I24), paroxysmal tachycardia (I47), atrial fibrillation and flutter (I48), hypertension (I10), and cardiac disease (I51). Figure 4 shows that using imaging phenotypes alone led to moderate AUC scores (e.g., 0.8361 and 0.8201 for myocardial infarction and ischemic heart diseases using with AdaBoost). Using the latent vector resulted in increased AUC scores (0.8557 and 0.8453). Combining both imaging phenotypes and the latent vector further improved the AUC scores (0.8762 and 0.8472), indicating the usefulness of the latent vector for cardiovascular disease classification.
@@
For the AdaBoost classifier, using feature sets comprising the latent vector, as well as the combination of phenotypes and the latent vector, consistently outperformed the performance of the phenotypes set alone (e.g., 0.8291 and 0.8316 for cardiac disease using latent vector and combined feature sets), implying that incorporating the latent vector improved the classification accuracy. The trend was particularly noticeable for myocardial infarction, hypertension, and cardiac diseases, where the combined phenotypes and latent vector feature set substantially improved the AUC scores (0.8762, 0.7738 and 0.8316 for myocardial infarction, hypertension and cardiac disease). The LDA and SVM classifiers demonstrated that of the three feature sets, the combined phenotypes and latent vector feature set achieved the highest AUC scores (e.g., 0.6728 and 0.6479 for hypertension with LDA and SVM). However, for certain diseases such as ischemic heart disease, classifiers using only phenotypes (e.g., 0.7381 and 0.7123 for ischemic heart diseases with LDA and SVM) outperformed those that used only the latent vector (0.7277 and 0.6975), but still fell short of their combination (0.7492 and 0.7214). Overall, the results show that integrating imaging phenotypes, the latent vector along with confounders provides the best discriminative feature set for classification.
@@
The latent delta has been shown to correlate with phenotypes such as LVM and LVEF (Figure 5(a)), which serve as indicators of cardiac structure and function. Conditions such as hypertension, lipid/cholesterol abnormalities, and diabetes can induce changes in these cardiac phenotypes. For example, hypertension likely results in an increased LVM and may be linked to a reduced LVEF due to the heart’s adaptation to prolonged high blood pressure. In a similar vein, diabetes can exert metabolic stress on the heart, which can lead to changes in cardiac volume and ejection fraction. These modifications in the structure and motion patterns of the heart, as captured by the latent delta, provide a mechanistic explanation for the associations observed in the PheWAS results. Consequently, significant correlations between the latent delta and various clinical outcomes highlight its potential as a valuable marker for linking cardiac phenotypes with broader health conditions.
@@ Discussion
Although this work advances the science in personalised cardiac modelling, there are several limitations. First, the personalised normative model relies on a restricted range of generating factors, including age, sex, weight, and height, as we aim to develop a standard healthy heart. Including additional elements in the future such as diseases or environmental factors such as air pollution and noise[42] could improve our understanding of their impacts on cardiac anatomy and function. Second, the model uses a cross-sectional dataset from the UK Biobank for both training and testing purposes. However, it does not include a benchmark for the progression of cardiac aging, which could be addressed by employing a longitudinal dataset to evaluate the model. Repeated scans are expected in the near future from the UK Biobank. Third, this study focuses on modelling the dynamic mesh sequence to describe cardiac shape and motion. It does not aim to model the underlying electrophysiology or biomechanics of the heart, which are also essential for cardiac modelling and understanding cardiac function [43, 44]. Additionally, the explainability of latent vectors could be explored, as understanding the specific information each latent dimension captures is crucial for clinical interpretation and validation.

### v2

@@ Abstract
Understanding the structure and motion of the heart is crucial for diagnosing and managing cardiovascular diseases, the leading cause of global death. There is wide variation in cardiac shape and motion patterns, that are influenced by demographic, anthropometric and disease factors. Unravelling the normal patterns of shape and motion, as well as understanding how each individual deviates from the norm, would facilitate accurate diagnosis and personalised treatment strategies. To this end, we developed a novel conditional generative model, MeshHeart, to learn the distribution of shape and motion patterns for the left and right ventricles of the heart. To model the high-dimensional and complex spatio-temporal mesh data, MeshHeart employs a geometric encoder to represent cardiac meshes in a latent space, followed by a temporal Transformer to model the motion dynamics of latent representations. Based on MeshHeart, we investigate the latent space of 3D+t cardiac mesh sequences and propose a novel distance metric termed latent delta, which quantifies the deviation of a real heart from its personalised normative pattern in the latent space. In experiments using a large cardiac magnetic resonance image dataset of 38,309 subjects from the UK Biobank, MeshHeart demonstrates a high performance in cardiac mesh sequence reconstruction and generation. Features defined in the latent space are highly discriminative for cardiac disease classification, whereas the latent delta exhibits strong correlations with clinical phenotypes in phenome-wide association studies. The code and the trained model will be released to benefit further research on digital heart modelling.
@@ Introduction
The heart is one of the most important and vital organs within the human body [1]. It is composed of four morphologically distinct chambers that function in a coordinated manner. The shape of the heart is governed by genetic and environmental factors [2, 3], as well as a remodelling process observed in response to myocardial infarction, pressure overload, and cardiac diseases [4, 5]. The motion of the heart follows a periodic non-linear pattern modulated by the underlying molecular, electrophysiological, and biophysical processes [6]. Unveiling the complex patterns of cardiac shape and motion will provide important insights for assessing the status of cardiac health in both clinical diagnosis and cardiovascular research [7, 8, 9, 10].
@@
The current state-of-the-art for assessing cardiac shape and motion is to perform analyses of cardiac images, e.g., cardiac magnetic resonance (MR) images, and extract imaging-derived phenotypes (IDPs) of cardiac chambers [11, 9]. Most imaging phenotypes, such as chamber volumes or ejection fractions, provide a global and simplistic measure of the complex 3D-temporal (3D+t) geometry of cardiac chambers [12, 11]. However, these global volumetric measures may not fully capture the dynamics and variations of cardiac function across individuals. Recent studies have shown that mesh-based cardiac shape and motion analyses can provide more detailed and clinically relevant insights [13, 14, 15, 16]. For example, Piras et al.[14] proposed to use spatio-temporal motion analysis to identify myocardial infarction. Gilbert et al.[15] highlighted stronger associations between cardiac risk factors and mesh-derived metrics in the UK Biobank dataset. Mauger et al.[16] showed that mesh-based motion metrics could independently predict adverse cardiac events. This underscores the importance of establishing a precise computational model of cardiac status to define what a normal heart looks like and moves like. Nevertheless, it is a non-trivial task to describe the normative pattern of the 3D shape or even 3D-t motion of heart, due to the complexity in representing high-dimensional spatio-temporal data.
@@
Recently, machine learning techniques have received increasing attention for cardiac shape and motion analysis [6, 17, 18]. Most existing research focuses on developing discriminative machine learning models, that is, training a model to perform classification tasks between different shapes or motion patterns [19, 20, 6, 8]. However, discriminative models offer only classification results and do not explicitly explain what the normative pattern of cardiac shape or motion looks like [21]. In contrast, generative machine learning models provide an alternative route. Generative models are capable of describing distributions of high-dimensional data, such as images [22, 23, 24], geometric shapes [25, 26, 27], or molecules [28, 29], which allow the representation of normative data patterns in the latent space of the model. In terms of generative modelling of the heart, recent developments focus on shape reconstruction and virtual population synthesis[13, 30, 31, 32, 33, 34]. For example, Xia et al. proposed a method that integrates statistical shape priors with deep learning for four-chamber cardiac shape reconstruction from images [35]. Gaggion et al. introduced HybridVNet, which combines CNNs with graph convolutions to perform shape reconstruction from multi-view images [36]. Dou et al. proposed a conditional flow-based variational autoencoder (VAE) for synthesising virtual populations of cardiac anatomy[37], and later developed a compositional generative model for multipart anatomical structures [38]. Beetz et al. introduced a variational mesh autoencoder that models population-wide variations in cardiac shapes with a hierarchical structure[39], and investigated the interpretability of the latent space extracted from a point cloud VAE [40]. Although generative models have been explored for cardiac shape reconstruction [36, 35], shape modeling [3, 39, 37, 38], image/video generation [41, 42, 43] and data augmentation [44], their application to personalised normative modeling of the heart from population data remains underexplored.
@@
Figure 1: An overview of the MeshHeart model. a, Model architecture: MeshHeart encodes a sequence of cardiac meshes using a mesh encoder $M_{\text{enc}}$ and encodes clinical factors using a conditional encoder $C_{\text{enc}}$. The encoder outputs across the time frames and along with distribution tokens $\mu_{token}$ and $\Sigma_{token}$, are processed by a temporal Transformer encoder $T_{\text{enc}}$. A Transformer decoder $T_{\text{dec}}$ and a mesh decoder $M_{\text{dec}}$ generate a 3D cardiac mesh sequence based on clinical factors. b, Given a set of clinical factors, an example of the generated mesh sequence across time frames. c, Conceptual framework: MeshHeart constructs a normative cardiac mesh sequence using personal information including age, sex, weight, and height. A real heart can be compared to its personalised norm by the latent vector. The latent delta $\Delta z$ is a distance defined between the latent vector of a synthetic normal heart (dark blue dot) and that of the real heart (red dot).
@@ Results
Utilising the latent representations learnt by MeshHeart, we assessed the ability of the model to generate new synthetic cardiac mesh sequences that mimic real heart dynamics. To evaluate the fidelity and diversity of the generation, we calculated the similarity between the distributions of real meshes and generated synthetic meshes. For each real heart in the test set ($n=4,000$), we applied MeshHeart to generate synthetic mesh sequences using the same clinical factors (age, sex, weight, and height) as the individual as the model input. During the generation stage, we chose 20 random samples from the Gaussian distribution of the latent space and generated the corresponding mesh sequences. For both real and synthetic meshes, clinically relevant metrics for cardiac structure and function were derived, including left ventricular end-diastolic volume (LVEDV), LV end-systolic volume (LVESV), LV ejection fraction (LVEF), LV myocardial mass (LVM), right ventricular end-diastolic volume (RVEDV), RV end-systolic volume (RVESV) and RV ejection fraction (RVEF). For each metric $m$, its probability distributions against age $P(m|c=age)$ and against sex $P(m|c=sex)$, were calculated. The similarity between real and synthetic probability distributions was quantified using the Kullback-Leibler (KL) divergence [48] and the Wasserstein distance (WD) [49], with a lower value denoting a higher similarity, i.e. better generation performance. KL divergence is a metric from information theory that evaluates the dissimilarity between two probability mass functions. Similarly, WD measures the dissimilarity between two probability distributions. MeshHeart’s ability to replicate real data distributions is quantitatively demonstrated in Figure 3(a). MeshHeart achieves lower KL and WD scores compared to existing methods, as shown by radar plots with the smallest area, suggesting that the synthetic data generated by the proposed model closely aligns with the real data distribution for clinically relevant metrics. Supplementary Tables S3 and S4 report the detailed KL divergence and WD scores for different methods.
@@
We also examined the latent representation learnt by MeshHeart using t-SNE (t-distributed Stochastic Neighbour Embedding) visualisation [50] as illustrated in Supplementary Figure S1. The t-SNE plot projects the 64-dimensional latent representation of a mesh, extracted from the last hidden layer of the Transformer Encoder $T_{\text{enc}}$, onto a two-dimensional space, with each point denoting a mesh. It shows 10 sample sequences. For each sample, the latent representations of the meshes across time frames form a circular pattern that resembles the rhythmic beating of the heart [51].
Figure 4: Comparison of disease classification performance, in terms of the AUC (area under the curve) scores, when different feature sets are used. The three feature sets being compared include "Phenotypes $+$ Confounders (Age, Sex, Weight, Height)", "Latent vector $+$ Confounders", and "Phenotypes $+$ Latent vector $+$ Confounders", with each feature set represented by a unique color in the plot. The three subplots show the performance of three different classifiers, AdaBoost, LDA, and SVM. The x-axis denotes the disease type and the y-axis denotes the AUC score. For the three feature sets, each feature set is compared to one of the other two features sets using the DeLong’s test. A single asterisk (*) denotes a notable difference ($p<0.05$), indicating that a feature set outperforms another feature set significantly, while a double asterisk (**) indicates that a feature set outperforms both of the other two feature sets significantly.
@@
After demonstrating the generative capability of MeshHeart, we explore its potential for clinical applications, in particular using its latent space which provides a low-dimensional representation of cardiac shape and motion. The latent feature analyses were conducted on 17,309 subjects. More than half (58.5%) had a reported diagnosis of at least one disease. We employ the latent vector $z$ of each mesh sequence, a 64-dimensional vector, as the feature for correlation analysis and for cardiac disease classification. Figure 5(a) shows that the latent vector exhibits strong correlations with conventional imaging phenotypes, such as LVM, LVEDV, and RVEDV etc. Figure 4 and Supplementary Table S6 compare the classification performance of six cardiac diseases when using different feature sets. The three evaluated feature sets include "Phenotypes + Confounders (Age, Sex, Weight, Height)", "Latent Vector + Confounders", and "Phenotypes + Latent Vector + Confounders". The classification performance is evaluated using the AUC (area under the curve) scores for three different classifiers: AdaBoost, linear discriminant analysis (LDA), and support vector machine (SVM). The six cardiovascular diseases include myocardial infarction (ICD-10 code I21), ischemic heart diseases (I24), paroxysmal tachycardia (I47), atrial fibrillation and flutter (I48), hypertension (I10), and cardiac disease (I51). Figure 4 shows that using imaging phenotypes alone led to moderate AUC scores (e.g., 0.8361 and 0.8201 for myocardial infarction and ischemic heart diseases using with AdaBoost). Using the latent vector resulted in increased AUC scores (0.8557 and 0.8453). Combining both imaging phenotypes and the latent vector further improved the AUC scores (0.8762 and 0.8472), indicating the usefulness of the latent vector for cardiovascular disease classification. These results demonstrate the model’s ability to discriminate not only between normal and abnormal cardiac states, but also among specific disease conditions.
@@
For the AdaBoost classifier, using feature sets comprising the latent vector, as well as the combination of phenotypes and the latent vector, consistently outperformed the performance of the phenotypes set alone (e.g., 0.8291 and 0.8316 for cardiac disease using latent vector and combined feature sets), implying that incorporating the latent vector improved the classification accuracy. The trend was particularly noticeable for myocardial infarction, hypertension, and cardiac diseases, where the combined phenotypes and latent vector feature set substantially improved the AUC scores (0.8762, 0.7738 and 0.8316 for myocardial infarction, hypertension and cardiac disease). While the model was trained using a normal healthy heart dataset, it has learnt a rich latent representation to encode diverse shape and motion patterns for different subpopulations in this large dataset. The resulting latent vector captures deviations in the latent space that are indicative of specific disease outcomes, as demonstrated by the experimental results. The LDA and SVM classifiers demonstrated that of the three feature sets, the combined phenotypes and latent vector feature set achieved the highest AUC scores (e.g., 0.6728 and 0.6479 for hypertension with LDA and SVM). However, for certain diseases such as ischemic heart disease, classifiers using only phenotypes (e.g., 0.7381 and 0.7123 for ischemic heart diseases with LDA and SVM) outperformed those that used only the latent vector (0.7277 and 0.6975), but still fell short of their combination (0.7492 and 0.7214). Overall, the results show that integrating imaging phenotypes, the latent vector along with confounders provides the best discriminative feature set for classification.
@@
The latent delta has been shown to correlate with phenotypes such as LVM and LVEF (Figure 5(a)), which serve as indicators of cardiac structure and function. Conditions such as hypertension, lipid/cholesterol abnormalities, and diabetes can induce changes in these cardiac phenotypes. For example, hypertension likely results in an increased LVM and may be linked to a reduced LVEF due to the heart’s adaptation to prolonged high blood pressure. In a similar vein, diabetes can exert metabolic stress on the heart, which can lead to changes in cardiac volume and ejection fraction. These modifications in the structure and motion patterns of the heart, as captured by the latent delta, provide a mechanistic explanation for the associations observed in the PheWAS results.
In Figure 5(b), the direction of effect shows the relationship between $\Delta z$ and the clinical outcome. A positive effect indicates that an increase in $\Delta z$ is associated with a higher probability of the outcome. In contrast, a negative effect indicates that a higher $\Delta z$ reduces the likelihood of the outcome. For example, a negative effect for birth trauma suggests that a higher $\Delta z$ is associated with a reduced likelihood of birth trauma. These directional effects provide insight into how deviations in cardiac structure and function relate to specific clinical outcomes, highlighting potential associations for further in-depth clinical investigation.
@@ Discussion
Although this work advances the science in personalised cardiac modelling, there are several limitations. First, the personalised normative model relies on a restricted range of generating factors, including age, sex, weight, and height, as we aim to develop a standard healthy heart. Including additional elements in the future such as diseases or environmental factors such as air pollution and noise[58] could improve our understanding of their impacts on cardiac anatomy and function. Second, the model uses a cross-sectional dataset from the UK Biobank for both training and testing purposes. However, it does not include a benchmark for the progression of cardiac aging, which could be addressed by employing a longitudinal dataset to evaluate the model. Repeated scans are expected in the near future from the UK Biobank. Third, this study focuses on modelling the dynamic mesh sequence to describe cardiac shape and motion. It does not aim to model the underlying electrophysiology or biomechanics of the heart, which are also essential for cardiac modelling and understanding cardiac function [59, 60]. Additionally, the explainability of latent vectors could be explored, as understanding the specific information each latent dimension captures is crucial for clinical interpretation and validation. Finally, our method does not incorporate long-axis images, which limits its ability to capture the mitral, tricuspid, or aortic valves for assessing valvular function. Mauger et al.[61] used 2-chamber and 4-chamber long-axis images to identify tricuspid and mitral valve points, so that the motion of the valve points can be tracked and modelled using principal component analysis.